OpenStudy (anonymous):

7sqrt{18}+8sqrt{50} 7sqrt{18}+8sqrt{50} @Mathematics

OpenStudy (anonymous):

\[7\sqrt{9\times2}+8\sqrt{2\times5\times5}\] thats as far as ive gotten

OpenStudy (anonymous):

\[61\sqrt{2}\]

OpenStudy (anonymous):

is that right so far?

OpenStudy (anonymous):

yeah it is right

OpenStudy (anonymous):

\[61\sqrt{2}\]

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

gr8 going tena

OpenStudy (anonymous):

thanks! its not easy lol